# 2]Centro de Matemática e Aplicações da Universidade da Beira Interior, Rua Marquês D’Ávila e Bolama 6200-001 Covilhã, Portugal

### v1

@@ Abstract
This work aims to explain the amplification of curvature perturbations on small scales arising from the phenomenon of self-resonance. By making a series expansion of the inflationary potential and employing perturbative techniques to solve for the inflaton field, we reformulate the Mukhanov-Sasaki equation into a Hill equation. Then, we derive expressions for the Floquet exponents corresponding to a potential having both cubic and quartic terms in its expansion. Our analytical results are then compared with numerical computations. Moreover, we propose potential applications including the generation of Primordial Black Holes, Scalar-Induced Gravitational Waves, and Oscillons.
@@ 1 Introduction
The cosmological models known as $\alpha$-attractors [1, 2, 3, 4] show an excellent agreement with Planck data [5] and have universal predictions of observables. For this reason, they have gained a lot of attention, as well as from the point of view of supergravity, where they are most naturally formulated. These models are characterized by the positive, real, dimensionless parameter $\alpha$. In this work we study the post-inflationary phase of preheating [6] for this class of models, characterized by the inflaton’s oscillation around its minimum. Specifically, our focus lies on the study of the amplification of curvature perturbations at small scales, driven by the phenomenon of self-resonance [7, 8], which arises when the potential deviates from quadratic behavior. While other works consider scenarios involving the coupling of the inflaton field to a daughter field, a gauge field, or a fermionic field, in this work we consider these couplings to be small, directing our attention towards the self-interactions of the field, where the inflaton drives its fluctuations. Self-resonance is mainly based on the Taylor series expansion of the potential and posterior transformation of the perturbation equation into a Hill equation, which gives us the Floquet’s exponents governing the amplification of the perturbations. Usually, the potential during this period is approximated by a parabola [6, 9, 10, 11], since the terms in the expansion of the potential higher than quadratic are not relevant. However, as $\alpha<1$, these terms become relevant, and thus self-resonance comes into play. The results obtained here can be extrapolated to any potential whose Taylor expansion is given by a quadratic part + higher order terms. Furthermore, the implications of our study extend to various phenomena. If the amplification from self-resonance reaches significant levels, perturbations could collapse and form Primordial Black Holes (PBHs) [12, 13, 14, 15], Scalar-Induced Gravitational Waves (SIGWs) [16, 17, 18, 19, 20], or coherent structures known as oscillons [21, 22, 23, 24, 25], we leave this as a future project [26, 27]. The contents of this work are organized as follows. We begin in Sec. 2 by revisiting the $\alpha$-attractor models. In Sec. 3 we study the effect of the self-resonance for a generic potential and compute the Floquet’s exponents. Then in Sec. 4 we show specifically for a T- and E-model the amplification of curvature perturbations and compare the results with a numerical computation. Sec. 5 focus on the importance of considering higher order terms in the expansion of the potential by comparing with the parabola approximation and finally in Sec. 6 we give some applications where this work can be useful. Conclusions are given in Sec. 7.
2 $\alpha$-attractor models
There are two types of $\alpha$-attractors, the so-called T- and E-models, whose potentials are given, respectively, by [4]
@@ 2 $\alpha$-attractor models
Here, $M$ is the inflaton mass, to be fixed by CMB normalization, $\phi$ is the scalar field and $\alpha$ is the parameter characterizing the models, with some constraints [^1] In [28] it is given a lower bound for the parameter $\alpha$, $\log_{10}\alpha=-4.2_{-8.6}^{+5.4}$ (95% C.L.), based on CMB constraints. We consider values of $\alpha$ above this lower bound for this work.. Expanding the potentials around $\phi=0$ we get
@@
This indicates that decreasing the parameter $\alpha$ enhances the importance of the cubic and quartic terms, even for relatively small values of $\phi$.
@@ 3.1 Perturbation theory for anharmonic oscillators
The background dynamics of the inflaton field $\phi$ are governed by the Klein-Gordon (KG) equation [29, 30]
$\ddot{\phi}+3H\dot{\phi}+\frac{\mathrm{d}V(\phi)}{\mathrm{d}\phi}=0\,,$ (5)
where a dot means derivation with respect to cosmic time, $H=\dot{a}/a$ is the Hubble rate, and $a$ is the scale factor of the universe. After inflation, the friction term is sub-dominant and thus we can neglect it. Later, we will revisit the issue of considering the Hubble expansion. Using perturbation theory [31] we can expand the field into powers of a small parameter $\epsilon$ as
$\phi=\epsilon\phi_{1}+\epsilon^{2}\phi_{2}+\epsilon^{3}\phi_{3}+\dots.$ (6)
If we now substitute this expansion into (5), use (4) and solve order by order we find that at order $\epsilon^{3}$ the solution behaves in a secular way since one of the driving terms in the differential equation has exactly the fundamental frequency of the system, i.e., $M$. To avoid this behavior we can use the Lindstedt-Poincaré renormalization procedure, where we expand both the field and the frequency, see [31] for further details. This expansion of the frequency is made by considering the new time variable [7, 8]
$\tau=t\sqrt{1-\epsilon^{2}}.$ (7)
Substitution into (5) of equations (4), (6) and changing to time variable $\tau$ gives the following expressions for the expansions of the field
$\phi_{1}=\Phi_{1}\cos(M\tau),$ (8a)
$\phi_{2}=\Phi_{2}(\cos(2M\tau)-3),$ (8b)
where
$\Phi_{1}=\frac{2M}{\sqrt{3\left|\lambda-\frac{10\lambda_{3}^{2}}{9M^{2}}\right|}},\qquad\Phi_{2}=\frac{\lambda_{3}}{6M^{2}}\Phi_{1}^{2}.$ (9)
The amplitude $\Phi_{1}$ is fixed when eliminating the secular term in the differential equation for $\phi_{3}$. The solution to this last is not shown since we will work up to order $\epsilon^{2}$. Regarding the asymmetric (with respect to the origin) term $\phi_{2}$ of the expansion, we see that it is proportional to $\lambda_{3}$. Thus we clearly see that the asymmetry of the potential is transferred to the solution of the field. For symmetric potentials ($\lambda_{3}=0\not=\lambda$), the solution to $\mathcal{O}(\epsilon^{2})$ is given by $\phi=\epsilon\phi_{1}+\mathcal{O}(\epsilon^{3})$, which is closer to the quadratic potential.
At this point, it is interesting to introduce the decay of the field. We have solved (5) without considering the Hubble friction term, which makes the field’s amplitude constant in time. To include it, it is enough to parameterize $\epsilon$ as
$\epsilon(\tau)=\epsilon_{\text{end}}\left(\frac{a_{\text{end}}}{a}\right)^{3/2},$ (10)
where $a_{\text{end}}$ is the scale factor evaluated at the end of inflation, $\tau_{\text{end}}$, and $\epsilon_{\text{end}}$ is obtained from (6) evaluated at $\tau=\tau_{\text{end}}$ up to the desired order. To simplify notation we omit the argument of $\epsilon$. Also, we neglect derivatives of $\epsilon$, since it is assumed to vary slowly.
@@ 3.2 Floquet theory for perturbations
The evolution of the scalar perturbations is controlled by the so-called MS variable $v_{\bm{k}}$ [30, 29, 9]. The MS equation expressed in the time variable $\tau$ is written as
$\tilde{v}_{\bm{k}}^{\prime\prime}+\left(\frac{k^{2}}{a^{2}}+\frac{\mathrm{d}^{2}V}{\mathrm{d}\phi^{2}}+3\phi^{\prime 2}-\frac{\phi^{\prime 4}}{2H^{2}}+\frac{3}{4}\left(\frac{\phi^{\prime 2}}{2}-V\right)+\frac{2\phi^{\prime}}{H}\frac{\mathrm{d}V}{\mathrm{d}\phi}\right)(1+\epsilon^{2})\tilde{v}_{\bm{k}}=0,$ (11)
where a tilde means derivation with respect to $\tau$, $\tilde{v}_{\bm{k}}=a^{1/2}v_{\bm{k}}$ is the re-scaled MS variable [32, 10, 11] and we have expanded $(1-\epsilon^{2})^{-1}\simeq 1+\epsilon^{2}+\mathcal{O}(\epsilon^{4})$, which comes from changing from $t$ to $\tau$ in $\tilde{v}_{\bm{k}}^{\prime\prime}$. Substituting in (11) the expansion of the potential (4) together with the expansion of the field (6) and making $z=M\tau=Mt\sqrt{1-\epsilon^{2}}$ we find, up to order $\epsilon^{2}$
$\frac{\mathrm{d}^{2}\tilde{v}_{\bm{k}}}{\mathrm{d}z^{2}}+\left(A_{k}+2q_{1}\cos(z)+2p_{1}\sin(z)+2q_{2}\cos(2z)+2p_{2}\sin(2z)+h.h.\right)\tilde{v}_{\bm{k}}=0,$ (12)
with
$A_{k}=\frac{k^{2}}{a^{2}M^{2}}+1+\left(1+\left(\frac{3}{2}+\frac{3\lambda}{2M^{2}}-\frac{\lambda_{3}^{2}}{M^{4}}\right)\Phi_{1}^{2}\right)\epsilon^{2},$ (13a)
$q_{1}=\frac{\lambda_{3}\Phi_{1}\epsilon}{M^{2}},$ (13b)
$p_{1}=-\left(\sqrt{\frac{3}{2}}-\sqrt{\frac{25}{6}}\right)\frac{\lambda_{3}\Phi_{1}^{2}\epsilon^{2}}{2M^{2}},$ (13c)
$q_{2}=\left(\frac{3\lambda}{2M^{2}}+\frac{\lambda_{3}^{2}}{3M^{4}}-\frac{3}{2}\right)\frac{\Phi_{1}^{2}\epsilon^{2}}{2},$ (13d)
$p_{2}=-\sqrt{\frac{3}{2}}\Phi_{1}\epsilon,$ (13e)
and where $h.h.$ in (12) stands for higher harmonics. We do not consider them since doing so would introduce a higher order in $\epsilon$ when calculating the Floquet exponent [33]. To derive (12) we have used also that the Hubble rate is given up to $\mathcal{O}(\epsilon^{2})$ by
$H^{2}=\frac{1}{3}\left(\frac{\dot{\phi}^{2}}{2}+V\right)=\frac{\epsilon^{2}M^{2}\Phi_{1}^{2}}{6}.$ (14)
Eqn. (12) is known as a Hill equation, a generalization of a Mathieu equation [7, 8, 9, 6]. Depending on the values of the parameters $A_{k}$, $q_{i}$, and $p_{i}$ the physical modes $(\frac{k}{a})^{-1}$ experience instability or stability as they evolve. Following Floquet’s theorem, the re-scaled MS variable evolves as $\tilde{v}_{\bm{k}}\sim e^{\int\mu_{k}(z)\mathrm{d}z}$, where $\mu_{k}$ are the so-called Floquet exponents. For $\Re(\mu_{k})>0$ we have exponential growth and the mode is said to be unstable. The Floquet exponents of the Hill equation are generally unknown but can be computed using the harmonic balance method. In this work, we will apply the method outlined in [7, 8] (see also Appendix A of [34]). In order to solve (11), the MS variable $\tilde{v}_{\bm{k}}$ is expanded into harmonics as
$\tilde{v}_{\bm{k}}=\sum_{n=0}^{\infty}e^{inz}C_{n}(z).$ (15)
@@
$\begin{split}\frac{\mathrm{d}C_{n}}{\mathrm{d}z}&=\frac{i}{2n}\left[(A_{k}-n^{2})C_{n}+q_{1}(C_{n+1}+C_{n-1})+ip_{1}(C_{n+1}-C_{n-1})+\right.\\ &\left.+q_{2}(C_{n+2}+C_{n-2})+ip_{2}(C_{n+2}-C_{n-2})\right],\end{split}$ (16)
where we are neglecting $\frac{\mathrm{d}^{2}C_{n}}{\mathrm{d}z^{2}}$ since we assume that $C_{n}$ are slowly varying. The first instability band emerges from examining the fundamental frequencies $n=\pm 1$ in (16). Tipically, to order $\epsilon^{2}$, the coefficients $C_{\pm 1}$ are coupled just one to each other. However, in our case, the presence of $q_{1}$ and $p_{1}$, along with $q_{1}$ and $p_{2}$ being of order $\epsilon$, the coefficients $C_{\pm 1}$ are also being coupled to $C_{0}$, $C_{\pm 2}$ and $C_{\pm 3}$ to order $\epsilon^{2}$. This additional coupling complicates the computation of the Floquet exponents. Nonetheless, we have found that it suffices for our purposes to consider the fundamental frequencies $n=\pm 1$ and their corresponding couplings with the nearest harmonics $n=0,2$ and $n=-2,0$, respectively, up to order $\epsilon^{2}$. By doing this, we derive the following differential matrix equation
$\frac{\mathrm{d}}{\mathrm{d}z}\begin{pmatrix}C_{1}\\ C_{-1}\end{pmatrix}\simeq\frac{i}{2}\begin{pmatrix}A_{k}-1-\frac{q_{1}^{2}+p_{1}^{2}}{A_{k}}-\frac{q_{1}^{2}+p_{1}^{2}}{A_{k}-4}&q_{2}-ip_{2}-\frac{(q_{1}-ip_{1})^{2}}{A_{k}}\\ -q_{2}-ip_{2}-\frac{(q_{1}+ip_{1})^{2}}{A_{k}}&1-A_{k}+\frac{q_{1}^{2}+p_{1}^{2}}{A_{k}}+\frac{q_{1}^{2}+p_{1}^{2}}{A_{k}-4}\end{pmatrix}\begin{pmatrix}C_{1}\\ C_{-1}\end{pmatrix}.$ (17)
Finally, remembering Floquet’s theorem, the re-scaled MS variable evolves as $\tilde{v}_{\bm{k}}\simeq e^{\int\mu_{k}(z)\mathrm{d}z}$ [9], where $\mu_{k}(z)$ are now given by the eigenvalues of the matrix in (17). Up to order $\epsilon^{2}$, these are given by [^2] Since we are interested only in amplification we will consider just the one with the $+$ sign.:
$\mu_{k}$ =±$=\pm\frac{1}{2}\biggl{[}q_{2}^{2}+p_{2}^{2}-(A_{k}-1)^{2}+2q_{1}^{2}\biggl{(}\frac{A_{k}-1}{A_{k}}+\frac{A_{k}-1}{A_{k}-4}-\frac{q_{1}^{2}}{2(A_{k}-4)^{2}}$ (18)
$-\frac{q_{1}^{2}}{A_{k}(A_{k}-4)}\biggr{)}+2p_{1}^{2}\left(\frac{A_{k}-1}{A_{k}}+\frac{A_{k}-1}{A_{k}-4}\right)-\frac{2q_{1}^{2}q_{2}}{A_{k}}-\frac{4p_{1}p_{2}q_{1}}{A_{k}}\biggr{]}^{1/2}.$
@@
where $\epsilon_{H}=-\dot{H}/H^{2}$ is the Hubble slow-roll parameter. From (19) we see that when the amplification of $\tilde{v}_{\bm{k}}$ exceeds $a^{3/2}$, the curvature perturbation will grow. Typically, for potentials close to parabolic, this amplification is as high as $a^{3/2}$, causing $\mathcal{R}_{\bm{k}}$ to remain approximately constant for the modes inside the resonance band [10]. However, as we demonstrate in the next section, deviating from quadratic behavior alters the scenario, producing a fast amplification in a specific range of modes.

### v2

# Self-resonance during preheating: The case of $\alpha$-attractor models

@@ Abstract
In this paper, for the first time, we obtain a new class of solutions for the Hill-type differential equations, which emerge in the preheating self-resonance phase of the expanding Universe. We study, in particular, the class of symmetric and asymmetric scalar field potentials coming from the so-called $\alpha$-attractor models of the early Universe cosmology. By making a series expansion of the potential and employing perturbative techniques we reformulate the Mukhanov-Sasaki equation, which captures the dynamics of the curvature perturbation in these models, into a Hill equation. This last includes higher-order terms that were never solved in the literature. Namely, those coming from the cubic and quartic contributions of the scalar field potential. Then, we derive the expressions for the Floquet exponents of the Mukhanov-Sasaki variable. Our analytical results are then compared with numerical computations, showing a good agreement and thus making this method valuable for obtaining theoretical predictions with new observational applications in the contexts of Primordial Black Holes and Scalar-Induced Gravitational Waves.
@@ 1 Introduction
The cosmological models known as $\alpha$-attractors Kallosh:2013daa ; Kallosh:2013hoa ; Kaiser:2013sna ; Kallosh:2015lwa show an excellent agreement with Planck data Planck:2018jri and have universal predictions of observables. For this reason, they have gained a lot of attention, as well as from the point of view of supergravity, where they are most naturally formulated. These models are characterized by the positive, real, dimensionless parameter $\alpha$. In this work, we study the post-inflationary phase of preheating for this class of models, characterized by the inflaton’s oscillation around its minimum. This has already been studied in the context of chaotic inflation from the relevant papers of Kofman, Linde, Starobinsky in the 90s Kofman:1997yn ; Kofman:1994rk . For the $\alpha$-attractors scenario Iarygina:2018kee ; Iarygina:2020dwe ; Krajewski:2018moi ; Lozanov:2017hjm , these oscillations produce a huge amplification of perturbations for small values of $\alpha$. This effect is called self-resonance and can be so strong that the non-linearity regime is reached in $\mathcal{O}(1)$ e-folds after the end of inflation. See B for details. The $\alpha$-attractor models have been also studied in the context of parametric resonance. However, while many works Iarygina:2018kee ; Iarygina:2020dwe ; Krajewski:2018moi ; Lozanov:2017hjm ; Jedamzik:2010dq ; Martin:2019nuw ; Sfakianakis:2018lzf ; Ballesteros:2024hhq ; Martin:2020fgl ; Shafi:2024jig ; Mahbub:2023faw ; Sang:2020kpd ; Zhang:2023hjk ; Child:2013ria focus on a quadratic minimum or rely on numerical simulations, we instead expand the potential up to fourth order and develop an analytical method to explain the huge amplification produced by the self-resonance effects. Specifically, and contrary to the usual procedure, our focus lies on the study of the amplification of curvature perturbations through the Mukhanov-Sasaki (MS) equation (see A). We consider these perturbations to be more relevant from the observational point of view and thus we find it necessary to develop the analytical tools necessary to explain its amplification during preheating. We perform a series expansion of the potential and posterior transformation of the MS equation into a Hill equation drazin_1992 ; nayfeh2008nonlinear , which gives us the Floquet’s exponents governing the amplification of the perturbations. This Hill equation, that frequently emerge in the context of preheating, has never been solved including the terms coming from both the cubic and quartic contributions of the expansion of the potential. The idea of including this terms in the expansion of the inflationary potential comes from the fact that, as $\alpha$ decreases, they become relevant, and thus self-resonance comes into play more robustly, as we will shortly see. There are two types of $\alpha$-attractors, the so-called T- and E-models, whose potentials are given, respectively, by Kallosh:2015lwa
@@
Here, $M$ is the inflaton mass, to be fixed by CMB normalization, $\phi$ is the scalar field and $\alpha$ is the parameter characterizing the models, with some constraints [^1] In Iacconi:2023mnw it is given a lower bound for the parameter $\alpha$, $\log_{10}\alpha=-4.2_{-8.6}^{+5.4}$ (95% C.L.), based on CMB constraints. We consider values of $\alpha$ above this lower bound for this work.. Expanding the potentials around $\phi=0$ we get
@@
This, as we pointed above, indicates that decreasing the parameter $\alpha$ enhances the importance of the cubic and quartic terms, even for relatively small values of $\phi$. The results obtained here can be extrapolated to any potential whose Taylor expansion is given by a quadratic part + higher order terms. Furthermore, the implications of our study extend to various phenomena. If the amplification from self-resonance reaches significant levels, perturbations could collapse and form Primordial Black Holes (PBHs) Zeldovich:1967lct ; Hawking:1971ei ; Carr:1974nx ; Carr:1975qj or Scalar-Induced Gravitational Waves (SIGWs) Maggiore:2007 ; Guzzetti:2016 ; Domenech:2021 ; Baumann:2007zm ; Ananda:2006af . We leave this as a future project draft-alpha-PBH ; draft-alpha-GWs .
The contents of this work are organized as follows. We begin in Sec. 2 by studying the effect of the self-resonance on a generic potential and computing the Floquet’s exponents. Then in Sec. 3 we show specifically for a T- and E-model the amplification of curvature perturbations and compare the results with a numerical computation. Sec. 4 focuses on the importance of considering higher order terms in the expansion of the potential by comparing with the parabola approximation and finally in Sec. 5 we give some applications where this work can be useful. Conclusions are given in Sec. 6, A shows the derivation of the MS-equation, and B shows the time until the non-linearity regime is reached as well as some comparisons with lattice simulations.
@@ 2.1 Perturbation theory for anharmonic oscillators
The background dynamics of the inflaton field $\phi$ are governed by the equation of motion of the field and Friedmann equations Baumann:2009ds ; Mukhanov:1990me
$\ddot{\phi}+3H\dot{\phi}+\frac{\mathrm{d}V(\phi)}{\mathrm{d}\phi}=0\,,$ (5a)
$H^{2}=\frac{1}{3M_{p}}\left(\frac{\dot{\phi}^{2}}{2}+V_{T,E}(\phi)\right),$ (5b)
respectively, where a dot means derivation with respect to cosmic time, $H=\dot{a}/a$ is the Hubble rate, $a$ is the scale factor of the universe, and $M_{p}$ is the Planck mass. After inflation, the friction term in (5a) is sub-dominant and thus we can neglect it. Later, we will revisit the issue of considering the Hubble expansion. Under this assumption, eqn. (5a) can be solved using perturbation theory drazin_1992 ; nayfeh2008nonlinear . Among all the perturbative methods available, we choose to work with the Lindsted-Poincaré Method [^2] Other methods include the Time Transformation Method BURTON1983543 , the Modified Lindsted-Poincaré Method CHEUNG1991367 , the Multiple Scales Method nayfeh2008nonlinear , or some methods based on Jacobian elliptic functions Kovacic-2016 . We however restrict to the Lindsted-Poincaré method due to its simplicity and good results.. First, we change the differential equation (5a) to a time domain where the solution is simple harmonic and of period $2\pi$. This is done by considering $\tau=\omega t$, where $\omega$ is the real frequency of oscillation of $\phi$. Due to the presence of higher order terms in the expansion of the potential, the frequency is not simply given by $\omega^{2}=M^{2}$. Instead, we parametrize it as $\omega^{2}=M^{2}(1-\beta^{2})$, where $\beta$ is a small, real parameter that carries the information about the departure of the system from quadratic behavior. We are assuming a negative sign in the parametrization of $\omega^{2}$ due to the negative sign in the parameters $\lambda_{T}$ and $\lambda_{E}$, which makes the potential to be, in general, wider than quadratic, thus decreasing the frequency of oscillations of the field. Next, the field is expanded into powers of the same small parameter $\beta$ as follows
$\phi=\beta\phi_{1}+\beta^{2}\phi_{2}+\beta^{3}\phi_{3}+\dots,$ (6)
with initial conditions $\phi_{i}(0)=\Phi_{i}$ and $\phi_{i}^{\prime}(0)=0$. If we now substitute this expansion into (5a) and use (4), we find that at the lowest order, $\mathcal{O}(\beta)$, the system behaves as a simple harmonic oscillator
$\phi_{1}=\Phi_{1}\cos\tau.$ (7)
Now, going to the next order, $\mathcal{O}(\beta^{2})$, we observe that $\phi_{2}$ is a combination of the zeroth, first and second harmonics, that is
$\phi_{2}=\frac{\lambda_{3}\Phi_{1}^{2}}{6M^{2}}\left(\cos(2\tau)-3+2\cos(\tau)\right)+\Phi_{2}\cos(\tau).$ (8)
Following the approach of the time transformation method CHEUNG1991367 ; BURTON1983543 ; nayfeh2008nonlinear , the initial conditions for $\phi_{i}$ ($i>1$) can be chosen arbitrarily, as long as the main initial condition $\phi(0)=\phi_{0}$ holds. Thus, we can choose the amplitude $\Phi_{2}$ to cancel out the first odd harmonic. In this sense, the second-order solution contains only even harmonics, which enable us to obtain a more accurate solution CHEUNG1991367 . Going now to the next order, $\mathcal{O}(\beta^{3})$, we encounter that in the differential equation, there are some terms proportional to $\cos(\tau)$ depending on $\lambda_{3}$ and $\lambda$, which produces a secular behavior drazin_1992 . This is because some of the driving terms in the differential equation have exactly the fundamental frequency of the system. To avoid this, we choose the following value for the amplitude $\Phi_{1}$
$\Phi_{1}=\frac{2M}{\sqrt{3\left|\lambda-\frac{10\lambda_{3}^{2}}{9M^{2}}\right|}},$ (9)
so that now the solution $\phi_{3}$ does not contain secular terms. Choosing again $\Phi_{3}$ so that it cancels out the first odd harmonic, we obtain a solution for $\phi_{3}$ in terms of just the third odd harmonic as
$\phi_{3}=\Phi_{1}^{3}\left(\frac{\lambda}{32M^{2}}+\frac{\lambda_{3}^{2}}{48M^{4}}\right)\cos(3\tau).$ (10)
We find it enough for our purposes to work up to order $\mathcal{O}(\beta^{2})$, for which the solution is
$\phi=\beta\Phi_{1}\cos(\tau)+\beta^{2}\frac{\lambda_{3}\Phi_{1}^{2}}{6M^{2}}(\cos(2\tau)-3).$ (11)
For the initial condition $\phi(0)=\phi_{0}$ to hold, we must solve for $\beta$ at $\tau=0$ which, as previously said, is the correction to the frequency, which we defined as $\omega^{2}=M^{2}(1-\beta^{2})$. Once solved it will be given in terms of the anharmonic terms of the potential, namely $\lambda_{3}$ and $\lambda$, and in terms of the amplitude of the field $\phi_{0}$. The larger the amplitude, the larger the correction to the frequency since the anharmonic terms became more dominant. For a small amplitude, we have $\omega^{2}\simeq M^{2}$ and thus we recover the quadratic case (parabolic potential). In this sense, we see that including the friction term, and thus the decay of the field, the correction term $\beta$ decreases and the system moves towards the quadratic approximation of the potential. Therefore, one can approximate this situation by choosing a $\beta$ that decays at the same rate as the field. That is
$\beta(\tau)=\beta_{0}\left(\frac{a_{0}}{a}\right)^{3/2},$ (12)
where $\beta_{0}$ is obtained from (11) at $\tau=0$, considered as the end of inflation, and $a$ is the scale factor of the universe, with $a_{0}=a(0)$. To simplify notation we omit the argument of $\beta$ and neglect its derivatives since it is assumed to vary slowly. Fig. 1 shows the decay of the normalized amplitude of the field for the T- and E-models with different values of $\alpha$ using the perturbative expansion (11). For comparison, it has also been included the amplitude of the numerical solution of the equation of motion of the field, eqn, (5a), which has been solved self-consistently with the Friedmann equation (5b). As one can observe, the agreement is good, but as we decrease the values of $\alpha$ the analytic and numerical solutions deviate during the first e-fold of preheating. This is due to the lack of higher order terms in the expansion (11), which contribute more as $\alpha$ decreases. However, as Fig. 2 shows, working up to second order in $\beta$ is enough for our purposes.
Figure 1: Amplitude of the field for E- and T-models with different values of alpha. In each case, the perturbative expansion (11) is shown in dotted green and the numerical solution is in continuous black.
The term $\beta^{2}\lambda_{3}\Phi_{1}^{2}/2M^{2}$ in the solution for $\phi_{2}$ is often called the drift or steady streaming term nayfeh2008nonlinear , characteristic of systems with even non-linearities ($\lambda_{3}\phi^{2}$). It carries information about the asymmetry of the potential, since as we observe it is directly proportional to $\lambda_{3}$. Thus we clearly see that the asymmetry of the potential is transferred to the solution of the field, which will have a significant impact on the self-resonance effect, as we will see. For symmetric potentials ($\lambda_{3}=0\not=\lambda$), this self-resonance effect is not as pronounced as in the asymmetric case, since the solution is given naturally in terms of odd harmonics, which is closer to the quadratic case.
@@ 2.2 Floquet theory for perturbations
The evolution of the scalar perturbations is controlled by the MS variable $v_{\bm{k}}$ Mukhanov:1990me ; Baumann:2009ds , where a suffix $\bm{k}$ means we are working in Fourier space. The MS equation expressed in the time variable $\tau$ can be written as
$\frac{\mathrm{d}^{2}\tilde{v}_{\bm{k}}}{\mathrm{d}\tau^{2}}+\left(A_{k}+2q_{1}\cos(\tau)+2p_{1}\sin(\tau)+2q_{2}\cos(2\tau)+2p_{2}\sin(2\tau)+h.h.\right)\tilde{v}_{\bm{k}}=0,$ (13)
where $h.h.$ stands for higher harmonics [^3] We do not consider higher harmonics since doing so would introduce a higher order in $\beta$ when calculating the Floquet exponent Amin:2010xe , and the coefficients $A_{k}$, $q_{i}$ and $p_{i}$, as well as the derivation of the MS-equation, are specified in A. Eqn. (13) is known as a Hill equation, a generalization of a Mathieu equation Hertzberg:2014iza ; Hertzberg:2014jza ; Jedamzik:2010dq ; Kofman:1997yn . Depending on the values of the parameters $A_{k}$, $q_{i}$, and $p_{i}$ the physical modes $(k/a)^{-1}$ experience instability or stability as they evolve. Following Floquet’s theorem, the re-scaled MS variable evolves as $\tilde{v}_{\bm{k}}\sim\exp{\bigl{[}\int\mu_{k}(\tau)\mathrm{d}\tau\bigr{]}}$, where $\mu_{k}$ are the so-called Floquet exponents. For $\Re(\mu_{k})>0$ we have exponential growth and the mode is said to be unstable. The Floquet exponents of the Hill equation are generally unknown but can be computed using the harmonic balance method nayfeh2008nonlinear , where the MS variable $\tilde{v}_{\bm{k}}$ is expanded into harmonics as
$\tilde{v}_{\bm{k}}=\sum_{n=0}^{\infty}e^{in\tau}C_{n}(\tau).$ (14)
@@
$C_{n}^{\prime}=\frac{i}{2n}\left[(A_{k}-n^{2})C_{n}+q_{1}(C_{n+1}+C_{n-1})+ip_{1}(C_{n+1}-C_{n-1})+q_{2}(C_{n+2}+C_{n-2})+ip_{2}(C_{n+2}-C_{n-2})\right],$ (15)
where we are neglecting the $C_{n}^{\prime\prime}$ since we assume that $C_{n}$ are slowly varying. The first instability band emerges from examining the fundamental frequencies $n=\pm 1$ in (15). Tipically, to order $\beta^{2}$, the coefficients $C_{\pm 1}$ are coupled just one to each other. However, in our case, the presence of $q_{1}$ and $p_{1}$, along with $q_{1}$ and $p_{2}$ being of order $\beta$, the coefficients $C_{\pm 1}$ are also being coupled to $C_{0}$, $C_{\pm 2}$ and $C_{\pm 3}$ to order $\beta^{2}$. This additional coupling complicates the computation of the Floquet exponents. Nonetheless, we have found that it suffices for our purposes to consider the fundamental frequencies $n=\pm 1$ and their corresponding couplings with the nearest harmonics $n=0,2$ and $n=-2,0$, respectively, up to order $\beta^{2}$. By doing this, we derive the following differential matrix equation
$\begin{pmatrix}C_{1}^{\prime}\\ C_{-1}^{\prime}\end{pmatrix}\simeq\frac{i}{2}\begin{pmatrix}A_{k}-1-\frac{q_{1}^{2}+p_{1}^{2}}{A_{k}}-\frac{q_{1}^{2}+p_{1}^{2}}{A_{k}-4}&q_{2}-ip_{2}-\frac{(q_{1}-ip_{1})^{2}}{A_{k}}\\ -q_{2}-ip_{2}-\frac{(q_{1}+ip_{1})^{2}}{A_{k}}&1-A_{k}+\frac{q_{1}^{2}+p_{1}^{2}}{A_{k}}+\frac{q_{1}^{2}+p_{1}^{2}}{A_{k}-4}\end{pmatrix}\begin{pmatrix}C_{1}\\ C_{-1}\end{pmatrix}.$ (16)
Finally, remembering Floquet’s theorem, the re-scaled MS variable evolves as $\tilde{v}_{\bm{k}}\simeq\exp{\bigl{[}\int\mu_{k}(\tau)\mathrm{d}\tau\bigr{]}}$ Jedamzik:2010dq , where $\mu_{k}(\tau)$ are now given by the eigenvalues of the matrix in (16). Up to order $\beta^{2}$, these are given by:
$\begin{split}\mu_{k}=\frac{1}{2}\Biggl{[}q_{2}^{2}+p_{2}^{2}-(A_{k}-1)^{2}+2q_{1}^{2}\Big{(}\frac{A_{k}-1}{A_{k}}+&\frac{A_{k}-1}{A_{k}-4}-\frac{q_{1}^{2}}{2(A_{k}-4)^{2}}-\\ &-\frac{q_{1}^{2}}{A_{k}(A_{k}-4)}\Bigr{)}+2p_{1}^{2}\left(\frac{A_{k}-1}{A_{k}}+\frac{A_{k}-1}{A_{k}-4}\right)-\frac{2q_{1}^{2}q_{2}}{A_{k}}-\frac{4p_{1}p_{2}q_{1}}{A_{k}}\Biggr{]}^{1/2}.\end{split}$ (17)
@@
where $\epsilon_{H}=-\dot{H}/H^{2}$ is the Hubble slow-roll parameter in cosmic time. Eqns. (17) and (18) give us the time-evolution of the curvature perturbation as a function of the anharmonic coefficients of the potential. This allows us to efficiently give an estimation of the amplification of perturbations along with the position of the peak in terms of the comoving wavenumber $k$. This has some applications as we show in Sec. 5. The Floquet’s exponents of the curvature perturbations are, to our knowledge, shown for the first time for an inflationary potential presenting both cubic and quartic terms in its expansion. This implies a great advantage from the numerical point of view since traditionally the codes designed to compute Floquet’s exponents must deal with highly oscillatory functions that increase the computational time and become rather inefficient. From (18) we see that when the amplification of $\tilde{v}_{\bm{k}}$ exceeds $a^{3/2}$, the curvature perturbation will grow. Typically, for potentials close to parabolic, this amplification is as high as $a^{3/2}$, causing $\mathcal{R}_{\bm{k}}$ to remain approximately constant for the modes inside the resonance band Martin:2019nuw . However, as we demonstrate in the next section, deviating from quadratic behavior alters the scenario, producing a fast amplification in a specific range of modes.